[Hallc_running] [New Logentry] RC Daily Update

wmhenry at jlab.org wmhenry at jlab.org
Thu Dec 12 15:20:03 EST 2019

Logentry Text:
After the RC meeting yesterday we tried to commission the Moller Polarimeter.  Unfortunetly the spot size at the target was large (~12mm).  We consulted JP and \
proceeded however when going from tune beam to CW the ion chambers were tripping with even the smallest amount of beam.  MCC contacted Jay B. and they decided t\
hey could rework the upstream optics in the morning.  We decided not to risk the target and backed out of the Moller procedure.  Further details can be found he\
re: (https://logbooks.jlab.org/entry/3751506)(https://logbooks.jlab.org/entry/3751531).  Towards the end of swing the chopper in the injector prevented the hall\
 from getting beam.  An expert was contacted and the issue was solved rather quickly ( < 1hr).  At the end of swing MCC was unable to send rastered beam due to \
a FSD.  If you want the long story it is here: (https://logbooks.jlab.org/entry/3752214).  The short story is we shoud call Mark J. when there is a raster issue\
 and the problem will be fixed with an access, hopefully sometime around the middle of swing tonight.  With Brad's advice, I decided to proceed through owl with\
 unrastered beam.  We did a BPM calibration (starting at 20uA) and took defocused SHMS and focused HMS data.  Some HMS data had the sieve slit in.  This morning\
 MCC tested their new optics tune for the Moller and the spot size looked much better (https://logbooks.jlab.org/entry/3752164).  Afterwards the ramp rate lock \
was implemented at MCC (https://logbooks.jlab.org/entry/3752262).  Currently Dave, Murchanna and I are commissioning the Moller.  We should finish sometime in t\
he middle of Swing.  If all goes well the Spin Dance with Hall A will happen tomorrow morning at ~8AM.  At some point tonight MCC wants the do a Load Test by se\
nding Hall A 150 uA at 1 pass and Hall C 30 uA at 5 pass.  We need to discuss what we want to do with that 30 uA of beam.  I suspect items #5 and #6 is the runp\
lan with the N2 reference cell (https://hallcweb.jlab.org/wiki/index.php/Polarized_Helium-3_Experiments).  Tonight (middle of swing??) there will be an opportun\
ity for an access to fix the raster and do some quick target polarimetry checks in the hall.

We should also discuss target work that could happen Saturday.  I talked to Walter and his techs are kind enough to come in again this Saturday if needed but th\
ey need to leave by 3pm.  They will be needed to rotate the spectrometers and open/seal up the target enclosure.  Dave Hamlett can also support the Radcon effor\
t on Saturday.  I told Walter I will let him know what we want to do tomorrow after the 8am meeting.  J.P. was coordinated with Steve L about the spectrometer r\

Other Issues

Large BPM fix

LASER RTD: We need to watch this RTD and stop the Laser if needed
https://logbooks.jlab.org/entry/3752058	 (Ethan's)

Target Polarimetry work:  I haven't taken the time to fully understand these log entries but it appears we did see some polarization in the pumping cell.



We will have our daily RC meeting 4:15pm, 2nd floor Counting House Meeting Room.

To join the Meeting via BlueJeans:
To join via Room System:
Video Conferencing System: bjn.vc -or-
Meeting ID : 861439920
To join via phone :
1) Dial:
+1.408.740.7256 (US (San Jose))
+1.888.240.2560 (US Toll Free))
+1.408.317.9253 (US (Primary, San Jose))
(see all numbers - http://bluejeans.com/numbers)
2) Enter Conference ID : 861439920


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3752336
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20191212/61f569df/attachment.html>

More information about the Hallc_running mailing list